# A Large-scale Empirical Study on Large Language Models for Election Prediction

### v1

@@ Abstract
Can Large Language Models (LLMs) accurately predict election outcomes? While LLMs have demonstrated impressive performance in healthcare, legal analysis, and creative applications, their capabilities in election forecasting remain uncertain. Notably, election prediction poses unique challenges: limited voter-level data, evolving political contexts, and the complexity of modeling human behavior. In the first part of this paper, we explore and introduce a multi-step reasoning framework for election prediction, which systematically integrates demographic, ideological, and time-sensitive factors. Validated on 2016 and 2020 real-world data and extensive synthetic personas, our approach adapts to changing political landscapes, reducing bias and significantly improving predictive accuracy. We further apply our pipeline to the 2024 U.S. presidential election, illustrating its ability to generalize beyond observed historical data. Beyond enhancing accuracy, the second part of the paper provides insights into the broader implications of LLM-based election forecasting. We identify potential political biases embedded in pretrained corpora, examine how demographic patterns can become exaggerated, and suggest strategies for mitigating these issues. Together, this project, a large-scale LLM empirical study, advances the accuracy of election predictions and establishes directions for more balanced, transparent, and context-aware modeling in political science research and practice.
A Large-scale Empirical Study on Large Language Models for Election Prediction
Chenxiao Yu1, Zhaotian Weng1, Yuangang Li1, Zheng Li2, Xiyang Hu3, Yue Zhao1 1University of Southern California  2Arima  3Carnegie Mellon University {cyu96374, wengzhao, yuangang, yzhao010}@usc.edu, winston@arimadata.com, xiyanghu@cmu.edu Corresponding authors
@@ 1 Introduction
Large Language Models (LLMs) have shown strong capabilities across diverse domains, including drug discovery Liu et al. (2024), legal analysis Chalkidis et al. (2022), and creative tasks Yang et al. (2022). Their strength lies in not only understanding and generating text but also leveraging vast common knowledge Roberts et al. (2020), simulating diverse personas Hu and Collier (2024), and modeling human behavior in complex social contexts Bommasani et al. (2021). LLMs excel in human-like reasoning Zhou et al. (2020); AlKhamissi et al. (2022) and have been applied to simulate decision-making in various settings Zhou et al. (2023); Ziems et al. (2024).
Notably, recent work has extended LLMs to political science, analyzing policy documents, campaign speeches, and public sentiment Xu (2022); Haq et al. (2023); Li et al. (2024b). Beyond language processing, their ability to integrate knowledge and apply human-like reasoning positions them as powerful tools for simulating the intricate dynamics of political decision-making Argyle et al. (2023); Bisbee et al. (2024).
Figure 1: Demonstration of three prompt designs in §2.2. V1 is the direct prompt on voter demographic information, while V2 introduces time-dependent information to capture candidates’ agenda and V3 also uses multi-step reasoning. In this example for 2020 Ohio result prediction, only V3 can accurately predict the results, demonstrating the importance of leveraging both time-dependent information and multi-step reasoning for election result prediction.
Background. Despite LLMs’ achievements in more straightforward political science tasks, their capacity to handle more challenging efforts, such as election prediction, is not well understood Lerer et al. (2022). LLMs’ access to extensive historical data suggests potential for forecasting, but election prediction introduces unique challenges. First, the high cost of acquiring voter-level data complicates both experimentation and model validation. Second, election forecasting involves modeling not only voter behavior, but also the shifting political context, making it unclear whether text-based data alone can capture the full complexity Graefe (2014). Third, accurate predictions require reasoning about multiple factorsHolbrook (2016). Whether LLMs can perform this more advanced reasoning remains unknown. Wei et al. (2022).
This Work. We present a large-scale empirical study to address two key questions in election prediction with LLMs: (i) how can LLMs be leveraged to achieve accurate election predictions, and (ii) beyond accuracy, what additional considerations are crucial for understanding LLM capabilities in election forecasting tasks?
Contribution 1: Multi-Step Reasoning Framework for Accurate Election Prediction (§2). To handle the scarcity of detailed voter-level data, we employ the Sync synthetic data generation framework Li et al. (2020b), which probabilistically reconstructs individual demographic and behavioral profiles from aggregated public datasets. We pair this synthetic data with real-world datasets, notably the American National Election Studies (ANES) Time Series Studies (2022), ensuring that our model aligns with actual voting behaviors. Our approach also adapts to evolving political conditions by incorporating time-dependent factors. We aggregate information from presidential campaign data, including candidates’ policy agendas and backgrounds, to align the model with changing political environments Holbrook (2016). Finally, we introduce a multi-step reasoning framework guided by Chain of Thought prompting Wei et al. (2022). By breaking down the prediction process into intermediate steps, the model can systematically integrate demographic details, ideological alignment, and time-sensitive information. This approach strengthens the model’s accuracy and mitigates biases and overfitting issues observed in simpler models. As shown in Figs. 1 and 2, we refine our pipeline through iterations. The final version significantly improves predictive accuracy and alignment with real-world data via extensive experiments.
Contribution 2: New Insights and Directions Beyond Accurate Prediction (§3). Our experiments uncover several key insights that extend beyond raw predictive accuracy. First, while our multi-step reasoning pipeline reduces bias and improves alignment with real-world outcomes, residual skewness persists, highlighting the influence of pretrained corpora on political predictions. Second, although LLMs can replicate known demographic voting patterns, they tend to exaggerate these differences, raising concerns about reinforcing stereotypes. Finally, our results indicate that integrating richer contextual information and refining prompt strategies can further enhance predictive reliability. We thus propose meaningful future directions on reducing political skewness in pretrained corpora and mitigating stereotype biases in election prediction.
@@
2 Part 1: Accurate Election Prediction via a Multi-Step LLM Pipeline
How can we effectively leverage LLMs to predict election results? In this work, we simulate each voter’s decision-making process by providing LLMs with detailed voter information and asking them to predict voter preferences. To achieve this, we focus on two key aspects: (1) establishing an evaluation framework with appropriate datasets that contain voter-level information, and (2) designing an LLM-based pipeline for accurate election predictions. In §2.1, we describe the datasets and the evaluation methods. We then provide an overview of our design approach in §2.2 with three progressive pipelines. Finally, we evaluate the performance of the pipelines on datasets in §2.3.
@@ 2 Part 1: Accurate Election Prediction via a Multi-Step LLM Pipeline
Figure 2: Progressive design of LLM pipelines for election predictions. V1: Demographic-only Prompting2.2.1) uses static demographic personas but lacks temporal context. V2: Time-based Prompting2.2.2) adds election-year policy shifts and candidate information, but the single-step prompt setup leads to overloaded inputs that reduce prediction accuracy. V3: Multi-step Reasoning2.2.3) divides the decision-making process into multiple steps, improving reasoning quality and producing results that closely match real-world outcomes. All versions aggregate individual predictions into state-level results to capture broader election trends.
@@ 2.1 Datasets, Evaluation, and Settings
Datasets. This study leverages two primary data sources: (1) Public Benchmarks: The American National Election Studies (ANES) 2016 and 2020 Time Series data Studies (2019, 2022), which provide comprehensive demographic and political ideology information for real respondents. This dataset is a benchmark to evaluate how accurately LLMs simulate voter-level behavior in alignment with real-world data. (2) For state-level voting simulations, we use voter-level synthetic data generated via advanced ML techniques based on aggregated information Li et al. (2020b). Personas are randomly sampled for each state at specified ratios, and predictions are compared with actual election outcomes from 2020 Federal Election Commission (2021) and 2024 NBC News (2024).
Both datasets contain non-personally identifiable [^1] This project has been reviewed by the IRB and considered exempt as the datasets are non-personally identifiable., voter-level information. Detailed partition and sampling methodologies are in Appx. B.1.
Evaluation Method. To evaluate performance on public benchmarks and state-level simulations, we firstly assess how closely LLM simulations align with actual voting results. The calculation follows the approach outlined in Xie et al. (2024); Argyle et al. (2023) and is defined as:
Predicted Proportion Ratio $P(s)$
$P(s)=\frac{\text{Republican Votes}}{\text{Republican Votes}+\text{Democratic Votes}}$ (1)
Here, $s$ represents the unit of analysis: it can refer to an individual persona (as in public benchmarks like ANES) or an entire state (as in state-level simulations). This ratio $P(s)$ measures the proportion of votes predicted for the Republican Party relative to the total votes for the two major parties, excluding those with no preference.
Hardware and LLM Settings. Our experiments were conducted on a GPU server with an AMD EPYC Milan 7763 processor with 6 NVIDIA RTX A6000 Ada GPUs. For the LLM, we use OpenAI’s GPT-4o model for election predictions. Meta’s LLaMA 3.1 (405B) model is employed in intermediate steps to provide neutral summarizations of time-dependent information, enhancing certain pipelines for capturing temporal dynamicsFeng et al. (2023).
@@ 2.2 Our Progressive Design of LLM Pipelines
In this section, we present our progressive design for generating voter-level election predictions using LLMs. As shown in Fig. 2, we develop three versions of the pipeline. Each version addresses a key shortcoming of its predecessor and integrates more detailed information and reasoning processes.
@@ 1. V1:
    Demographic-only Prompting (§2.2.1): This initial version uses static demographic personas to guide voter-level predictions. It is simple but does not capture changes in candidates’ focus over time.
@@ 2. V2:
    Single-step Prompting with Time-based Information (§2.2.2): Here, we add election-year-specific details, such as policy agendas and candidate backgrounds, to make predictions more dynamic. However, packing all information into a single prompt can overwhelm the model, limiting its reasoning depth.
@@ 3. V3:
    Multi-step Reasoning with Contextual Information (§2.2.3): This version separates the prediction process into a sequence of steps. By breaking down the reasoning, the model can integrate demographics, candidate profiles, and political context better, for predictions that better match real-world results.
2.2.1 Version 1: Demographic-only Prompting
This initial version directly prompts the LLM with a persona that includes demographic attributes (e.g., age, gender, income) and asks how that persona would vote Xie et al. (2024); Argyle et al. (2023).
@@ 2.2.1 Version 1: Demographic-only Prompting
Specifying the year as 2020 avoids confusion because the LLM’s knowledge ends in 2023. The listed voting options follow Pew Research Center’s 2014 Political Polarization and Typology Survey Pew Research Center (2014).
Limitations: This version cannot adapt predictions to different time periods. Without incorporating temporal changes in candidates’ agendas or public opinion, predictions remain static, limiting their usefulness for evolving electoral contexts.
2.2.2 Version 2: Single-step Prompting with Time-based Information
Accurately modeling elections requires accounting for macro-level factors and time-specific variations Gao et al. (2022b). To improve realism, we extend our pipeline by incorporating election-year data from Ballotpedia, a widely used platform that provides campaign agendas, key policy positions, and candidates’ biographical details.
Ensuring that this time-based information is conveyed neutrally is important, given documented political biases in LLMs Feng et al. (2023). We compared GPT-4o and LLaMA3-405B for summarizing these details, and found that LLaMA3-405B provided more balanced outputs. We then integrated these balanced summaries into the prompts:
@@ 2.2.2 Version 2: Single-step Prompting with Time-based Information
Limitations:. Although adding time-dependent context makes predictions more dynamic, it introduced a skew towards Democratic candidates in both “deep red” states and key swing states. For example, predictions across five historically Republican states (e.g., Alabama, South Carolina) and multiple swing states (e.g., Texas, Florida) leaned more Democratic than expected. This pattern aligns with earlier observations that GPT-4o leans liberal Feng et al. (2023). Interestingly, when tested on the ANES 2020 dataset Studies (2022), which includes ideological self-placement, the skew was less pronounced. Additional analysis showed that removing ideological self-placement shifted outcomes significantly, suggesting that such ideological cues help correct biases introduced by purely demographic or time-based prompts.
2.2.3 Version 3: Multi-step Reasoning with Contextual Information
To address the limitations in Version 2, we introduce a multi-step prompting pipeline. Drawing on the Chain of Thought prompting strategy Wei et al. (2022), this approach decomposes the prediction process into intermediate steps, enabling more systematic reasoning and improved accuracy.
The procedure involves two main steps: (1) Conservative-Liberal Spectrum Placement: The model is given a persona and current policy positions of both parties, and is asked to position the persona on a conservative-liberal spectrum. (2) Extended Persona and Voting Simulation: This ideological self-placement is then incorporated into the persona. Together with time-based information, this enriched context is used in the second step to simulate voting behavior.

### v3

# A Large-Scale Simulation on Large Language Models for Decision-Making in Political Science

@@ Abstract
While LLMs have demonstrated remarkable capabilities in text generation and reasoning, their ability to simulate human decision-making—particularly in political contexts—remains an open question. However, modeling voter behavior presents unique challenges due to limited voter-level data, evolving political landscapes, and the complexity of human reasoning. In this study, we develop a theory-driven, multi-step reasoning framework that integrates demographic, temporal and ideological factors to simulate voter decision-making at scale. Using synthetic personas calibrated to real-world voter data, we conduct large-scale simulations of recent U.S. presidential elections. Our method significantly improves simulation accuracy while mitigating model biases. We examine its robustness by comparing performance across different LLMs. We further investigate the challenges and constraints that arise from LLM-based political simulations. Our work provides both a scalable framework for modeling political decision-making behavior and insights into the promise and limitations of using LLMs in political science research.
A Large-Scale Simulation on Large Language Models for Decision-Making in Political Science
Chenxiao Yu1, Jinyi Ye1, Yuangang Li1, Zhaotian Weng1, Zheng Li2, Emilio Ferrara1, Xiyang Hu3, Yue Zhao1 1University of Southern California  2Arima  3Arizona State University {cyu96374, jinyiy, yuangang, wengzhao, emiliofe, yzhao010}@usc.edu winston@arimadata.com, xiyanghu@asu.edu Corresponding authors
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong capabilities in processing and generating text, drawing on vast amounts of knowledge to assist with tasks in fields like scientific discovery Liu et al. (2024), law Chalkidis et al. (2022), and creative work Yang et al. (2022). Beyond text generation, they show emerging reasoning abilities that allow them to approximate human-like thought processes Zhou et al. (2020); AlKhamissi et al. (2022) and model human behavior Bommasani et al. (2021). However, LLMs still struggle to capture the deeper psychological and social mechanisms that drive human decision-making, making their simulations less reliable in real-world contexts Zhou et al. (2024). To address this, researchers have started incorporating insights from social science into LLM-based models. Recent studies have explored how LLMs can simulate economic decision-making Ross et al. (2024), public opinion dynamics Chuang et al. (2024), and social-psychological mechanisms like collaboration and conformity Zhang et al. (2024a).
Despite these advances, the application of LLMs to political decision-making remains underexplored. Voting behavior is one of the most fundamental decision-making processes in political science. LLMs are well-suited for this task because they have shown strong zero- and few-shot capabilities in simulating human dynamics, like political homophily in social networks Chang et al. (2024). Just as they have been used to estimate politicians’ ideological positions Wu et al. (2023), they could also approximate the average voter’s behavior, providing a scalable way to analyze political preferences at a broader level. Election simulations naturally emerge as a structured application of this approach. Unlike abstract ideological simulations, election outcomes offer a clear ground truth—real-world state- and county-level election data—making election simulation an ideal testbed for evaluating LLMs’ reasoning and predictive abilities in political science. If successful, this approach could extend to downstream applications, such as forecasting public reactions to policy changes, where traditional large-scale surveys and experiments are often costly and time-consuming.
Yet, accurately simulating voting behavior presents a number of challenges. First, LLMs inherit political biases from the data they are trained on, which can skew their predictions in politically sensitive tasks Feng et al. (2023). Second, voting decision-making is shaped by various factors, including demographics, location, ideology, and party affiliation Levendusky (2009); Abramowitz and Saunders (2008), but the high cost of acquiring voter-level data complicates both experimentation and model validation. Third, a large-scale election simulation should account not only for individual voter behavior but also for the shifting political context, but it remains unclear whether text-based data alone is sufficient to capture these information Graefe (2014). Fourth, accurate simulations may require multi-step reasoning Holbrook (2016), yet how to effectively integrate political science insights into LLMs’ reasoning processes—and whether LLMs can handle this level of complexity—remains an open question Wei et al. (2022).
This Work. We present a large-scale simulation study exploring how LLMs can model human decision-making in political science, focusing on voter behavior in U.S. elections. We develop a theory-driven, multi-step reasoning framework that incorporates demographic, ideological, and temporal factors to model political decision-making at scale. We evaluate our framework on different LLMs, compare their robustness and predictive performance, and investigate biases and limitations that emerge in large-scale political simulations. Our study addresses three key research questions:
RQ1: How can LLMs be used to simulate human decision-making in political science?
RQ2: How do different LLMs perform, and how robust are their election simulations?
RQ3: What limitations arise when using LLMs to model political decision-making?
@@
Contribution 1: A Theory-Driven Multi-Step Reasoning Pipeline for Accurate Election Simulation (§2).
We propose a theory-driven, multi-step reasoning pipeline to simulate voter decision-making, incorporating demographic, temporal, and ideological factors. To address the lack of detailed voter-level data, we use the Sync synthetic data generation framework Li et al. (2020b), which probabilistically reconstructs individual demographic and behavioral profiles from aggregated public datasets. We then align the personas with real-world voter data from American National Election Studies (ANES)  [^1] https://electionstudies.org/. Our approach also adapts to evolving political conditions by integrating temporal factors, such as candidates’ policy agendas and backgrounds Holbrook (2016).
Furthermore, building on political science studies on ideological sorting—the process by which voters increasingly align their political ideology with their party affiliation over time Levendusky (2009), we introduce ideology inference as an intermediate reasoning step. Using Chain-of-Thought prompting Wei et al. (2022), our model first predicts ideology based on demographics and behavioral data, which then influences party affiliation and voting preferences.
As shown in Fig 1, we refine our pipeline iteratively, incorporating demographics, political context, and ideological inference at each step. The final model significantly improves in simulation accuracy and alignment with real-world election.
Contribution 2: Challenges and Limitations in Large-Scale Political Simulations (§3). Our analysis reveals three important challenges in LLM-based political simulations. First, LLMs inherit systematic biases from their pretraining data, leading to a persistent left-leaning skew in simpler pipelines, with multi-step reasoning reducing but not eliminating this bias. Second, LLMs exaggerate demographic voting patterns, amplifying stereotypes related to gender, race, and education. Third, LLMs overestimate the influence of ideology on voting behavior, producing higher-than-real-world correlations between ideology and voting preference, a phenomenon referred to in previous work as “hyper-accuracy distortion” Aher et al. (2023). By these findings, we propose future research directions focused on debiasing training data, refining demographic calibration, and introducing human-in-the-loop techniques to improve the accuracy and reliability of LLM-based political simulations.
2 Accurate Simulation of Human Voting Behavior via a Multi-Step LLM Pipeline (RQ1, 2)
@@
Figure 1: Progressive design of LLM pipelines for voter simulation. V1: Demographic-Only Prompting2.2.1) uses static personas but lacks temporal context. V2: Time-Based Prompting2.2.2) adds election-year data V3: Multi-Step Reasoning2.2.3) structures decision-making into steps, improving reasoning and alignment.
How can we use LLMs to simulate human voting behavior in political science? In this work, we simulate each voter’s decision-making process by providing LLMs with detailed voter information and asking them to predict voter preferences.
To achieve this, we focus on two key components: (1) building a robust evaluation framework using datasets that contain voter-level information, and (2) designing a theory-driven Bafumi and Shapiro (2009); Pew Research Center (2014) LLM-based pipeline for accurate election simulation.
In §2.1, we present the datasets and evaluation methods. Next, we outline our design approach in §2.2, introducing three progressive pipelines, incorporating demographics, political context, and ideological inference at each step. Finally, we evaluate these pipelines by comparing their predictions with real-world outcomes in §2.3.
@@ 2.1 Datasets, Evaluation, and Settings
Datasets. This study leverages two primary data sources: (1) Public Benchmarks: The American National Election Studies (ANES) 2016 and 2020 Time Series data Studies (2019, 2022), which provide detailed demographic, ideology, and party affiliation information from real respondents. This dataset serves as a benchmark to evaluate how well LLMs simulate voter-level behavior in alignment with real-world patterns. (2) Large-Scale State-level Synthetic Voter Persona Dataset: A dataset of over 330,000 synthetic personas, generated using advanced ML techniques based on aggregated population census data and commercial datasets Li et al. (2020b). Personas are randomly sampled for each state at specified ratios, and their predicted voting outcomes are compared to actual U.S. election results from 2020 Federal Election Commission (2021) and 2024 NBC News (2024). Both datasets contain non-personally identifiable voter-level information [^2] This project has been reviewed by the IRB and exempt, as the datasets do not include personally identifiable info.. Detailed partitioning and sampling methodologies are provided in Appx. A.1.
Evaluation Method. To evaluate performance on public benchmarks and state-level simulations, we assess how closely LLM simulations align with actual voting results. The calculation follows the approach outlined in Argyle et al. (2023):
Predicted Voting Ratio $P(s)$
$P(s)=\frac{\textit{Republican Votes}}{\textit{Republican Votes}+\textit{Democratic Votes}}$ (1)
Here, $s$ represents the unit of analysis, which can refer to cross-regional samples (e.g., public benchmarks like ANES) or an entire state (e.g., state-level simulations). The ratio $P(s)$ measures the number of votes predicted for the Republican Party relative to the total votes for the two major parties, excluding those who express no preference.
LLMs and Hardware Settings. Our experiments utilized OpenAI’s GPT-4o and Meta’s LLaMA 3.1-70B model for the primary simulations. Meta’s LLaMA 3.1 (405B) model was employed in intermediate steps to provide neutral summarizations of time-dependent information Feng et al. (2023). Furthermore, we tested Qwen-72B and DeepSeek-V3 to measure systematic differences between models. For the hardware setup, we employed six NVIDIA RTX A6000 Ada GPUs and an 8-way NVIDIA A100 GPU cluster, with AMD Milan processors to execute tasks across different models.
@@ 2.2 Our Progressive Design of LLM Pipelines
In this section, we present our progressive design for generating voter-level behavior simulation using LLMs. As shown in Fig. 1, we develop three versions of the pipeline. Each version addresses a key shortcoming of its predecessor and integrates more detailed information and reasoning processes.
@@ 1. V1:
    Demographic-Only Prompting (§2.2.1): This baseline approach uses static demographic personas for voter-level simulations. While straightforward, it does not account for shifts in presidential candidates’ policy priorities over time.
@@ 2. V2:
    Single-Step Prompting with Time-Sensitive Information (§2.2.2): Here, we add election-year-specific details, like policy agendas and candidate backgrounds. However, packing all information into a single prompt may overwhelm the model, limiting reasoning depth.
@@ 3. V3:
    Multi-Step Reasoning with Ideology Inference (§2.2.3): This version structures the simulation into sequential steps, allowing the model to better integrate demographics, political ideology, and political context for more accurate real-world predictions.
2.2.1 V1: Demographic-Only Prompting
This initial version prompts the LLM with a persona’s demographics (e.g., age, gender, income) to simulate voting behavior Argyle et al. (2023). To prevent confusion, we specify the year as 2020, ensuring alignment with the LLM’s training data, which extends through 2023. The listed voting options follow Pew Research Center’s 2014 Political Polarization and Typology Survey Pew Research Center (2014).
@@ 2.2.1 V1: Demographic-Only Prompting
Please answer the following question as if you were the resident:
1. As of today, will you vote for the Democratic Party (Joe Biden), the Republican Party (Donald Trump), or do you have no preference?
@@
Limitations: This version lacks adaptability to different election cycles. Without accounting for shifts in candidate agendas or public opinion, its predictions remain static, limiting relevance in changing electoral contexts.
2.2.2 V2: Single-Step Prompting with Time-Sensitive Information
Accurately modeling elections requires accounting for macro-level factors and time-specific variations Gao et al. (2022). To improve realism, we extend our pipeline by incorporating election-year data from Ballotpedia [^3] https://ballotpedia.org/Main_Page, a widely used platform that provides campaign agendas, key policy positions, and candidate biographies. Given the documented political biases in LLMs Feng et al. (2023), ensuring that this time-based information is conveyed neutrally is crucial. We compared GPT-4o and LLaMA3-405B for summarizing these details and found that LLaMA3-405B produced more balanced outputs. These refined summaries were then integrated into the prompts.
@@ 2.2.2 V2: Single-Step Prompting with Time-Sensitive Information
Limitations: While incorporating time-dependent context makes predictions more dynamic, it does not eliminate inherent political biases Feng et al. (2023), which can still distort simulations of human behavior (see §2.3).
2.2.3 V3: Multi-Step Reasoning with Ideology Inference
Domain theory-driven design has been demonstrated to significantly improve the performance of LLMs in modeling the human decision-making process Chuang et al. (2024); Xie et al. (2024). Over the past few decades, political ideology has become increasingly aligned with party affiliation and partisanship Bafumi and Shapiro (2009), as well as with policy preferences and voting behavior in the United States Pew Research Center (2014); Levendusky (2009); Abramowitz and Saunders (2008) and in global political contexts Bornschier et al. (2021).
Building on these insights, we introduce a multi-step prompting pipeline inspired by Chain of Thought prompting Wei et al. (2022). This approach decomposes the prediction process into structured steps, enhancing reasoning and improving accuracy. The method consists of two key stages: (1) Political Ideology Inference: The model receives a persona along with current party policy positions and determines where the persona falls on the conservative-liberal spectrum. (2) Extended Persona and Voting Simulation: The inferred ideology is integrated into the persona, combined with time-based contextual information, and used to simulate voting behavior.